# Weingarten calculus with virtual isometries

### v2

@@ Abstract.
In this paper, we develop a novel approach to the Weingarten calculus employing the notion of virtual isometries. Traditionally, Weingarten calculus provides explicit formulas for integrating polynomial functions over compact matrix groups with respect to the Haar measure, yet faces limitations when evaluating high-degree integrals due to the non-invertibility of the associated matrices. We revisit these classical computations from a new perspective: by constructing Haar-distributed matrices via sequences of complex reflections, we derive new recursive structures for the Weingarten functions across different dimensions. This framework leads to two main results: (1) an explicit Weingarten calculus for complex reflections, yielding systematic moment computations for associated rank-one matrices, and (2) a novel convolution formula that connects Weingarten functions in dimension $n$ to those in dimension n1n-1 through the introduction of ascension functions in the symmetric group algebra. Our approach not only provides a unified treatment for unitary groups, but also sheds light on the algebraic and probabilistic aspects of high-degree integral computations. Several examples and applications are presented.
@@ 1.1. Weingarten calculus
An important case – on which we focus in this paper – is the unitary group $\mathrm{U}(n)$, which consists of all $n\times n$ unitary matrices. This case is important because every compact matrix group is conjugate to a subgroup of the unitary group. Let $U=(u_{ij})_{1\leq i,j\leq n}$ be a Haar-distributed matrix from $\mathrm{U}(n)$. From the properties of the Haar measure, when $k$ and $l$ are different nonnegative integers, the expectation of the following form is zero:
@@
for $i_{p},j_{p},i_{p}^{\prime},j_{p}^{\prime}\in\{1,2,\dots,n\}$ and $p\geq 1$. According to the Weingarten formula [Col03], when $k=l$, this moment can be expressed in the form of a sum over the symmetric group $S_{k}$:
@@
Here $\delta(p,q)$ is Kronecker’s delta, and $\operatorname{Wg}_{k,n}$ is a class function on $S_{k}$ known as the (unitary) Weingarten function. See also [Sam80]. The Weingarten function has various representations, but when $k\leq n$, we can immediately obtain the representation
@@
The Weingarten function has been the subject of extensive research. This includes expansions using characters of symmetric groups, expressions through Jucys–Murphy elements, and series expansions as generating functions for monotone factorizations of permutations [Nov10, MN13]. In addition, a relation identity has been obtained between the Weingarten functions on $S_{k}$ and those on $S_{k-1}$ [CM17].
@@ 1.2. The problem of high degree integrals
A common shortoming of Weingarten calculus is that if one wants to evaluate an integral of degree strictly more than $n$ on the unitary or orthogonal group of dimension $n$, the Weingarten matrix is not invertible anymore. For details, we refer, for example, to [CMN22]. There exist fixes to this problem. The first one is to take the pseudo-inverse of the Weingarten matrix when it is not invertible. It is important to note that while a Weingarten formula exists irrespective of the degree, the Weingarten function is not unique when the degree becomes larger than the dimension.
We are aware of two additional routes to compute polynomial integrals of high degree against the Haar measure, and describe them briefly in what follows. Both workarounds are successful from a theoretical point of view in the sense that they allow to compute in principle the integrals. However, at this point we are unable to extract any useful information for concrete applications. The first attempt can be found as a remark of a paper of Collins and Śniady [CŚ06], and the authors, along with Fukuda, have a joint working project to expand on this topic, [CFM25]. Roughly, the idea is that any formal computation in the dimension, while in principle only well defined for a dimension that is large enough – and potentially larger than the dimension in which one wants to compute, doing the calculation formally, simplifying, and then evaluating in the target dimension will systematically give the correct result.
The second route was taken by Cioppa and Collins in the following unpbulished paper [CC13]. Here, the idea is to replace the indicator functions which are simple to compute but do not yield an orthogonal Gram matrix by much more complicated functions obtained from the Gelfand-Zetlin basis, i.e. exhibit matrix elements of the commutant of tensor representations of the unitary group.
The third approach is novel – it is actually the motivation for this paper. It comes from the following simple idea: from the point of view of differential geometry and measure theory, $\mathrm{U}(n)/\mathrm{U}(n-1)=S^{2n-1}$, and thus $\mathrm{U}(n)=S^{2n-1}\times\mathrm{U}(n-1)$ and this extends to the level of product of uniform measures. For a rigorous statement, we refer to Lemma 2.1. Therefore, we can understand the Haar measure at the level $n$ from the Haar measure at the level n1n-1 and the uniform measure on the sphere. But it turns out that the uniform measure on the sphere can be understood with arbitrary order directly thanks to a Gaussianization trick. This has been used in [BC08], and we also refer to the remark after Lemma 2.2 for more details.
@@ 1.3. Virtual isometries
Due to [BNN13] (see also [MNN20]), virtual isometries can be constructed inductively as follows: We equip $\mathbb{C}^{n}$ with the canonical scalar product: $\langle\bm{v},\bm{w}\rangle=\sum_{j=1}^{n}v_{j}\overline{w_{j}}$ for $\bm{v}=(v_{1},\dots,v_{n})^{\mathtt{T}}$, $\bm{w}=(w_{1},\dots,w_{n})^{\mathtt{T}}\in\mathbb{C}^{n}$. Here $A^{\mathtt{T}}$ and $\bm{v}^{\mathtt{T}}$ stand for the transpose of a matrix $A$ or a vector $\bm{v}$. Let $\bm{e}_{1},\dots,\bm{e}_{n}$ be the standard orthonormal basis of $\mathbb{C}^{n}$. Consider a sequence $(\bm{x}_{n})_{n\geq 1}$ of column vectors, where each $\bm{x}_{n}$ lies on the complex unit sphere of $\mathbb{C}^{n}$ for $n\geq 1$. We define a sequence of unitary matrices $r_{n}\in\mathrm{U}(n)$ in the following way:
@@
The virtual isometry is consistent with Haar measure. Indeed, if $(X_{n})_{n\geq 1}$ is a sequence of independent random vectors, where each $X_{n}$ is uniformly distributed on the complex unit sphere of $\mathbb{C}^{n}$, and if we let $R_{n}$ and $U_{n}$ be the corresponding random matrices $r_{n}$ and $g_{n}$, respectively, in the above constructions, then $U_{n}$ follows the Haar measure on $\mathrm{U}(n)$.
A similar construction method for Haar measure is also provided by A. Hurwitz, refer to [DF17]. Virtual isometries have proven to be valuable in the analytic aspects of random matrix theory, enabling the attainment of almost sure results due to the correlations they provide between sequences of Haar-distributed unitary matrices, refer to [Ass22, CNN17].
@@ 1.4. Main results
The first main result is Theorem 2.4, where focus on a random complex reflection $R_{n}$ used in a construction involving virtual isometries. We establish Weingarten calculus for $R_{n}$, which provides a systematic method for computing the moments of the matrix elements of $R_{n}$. Instead of working directly with $R_{n}$, we will deal with $P_{n}:=I_{n}-R_{n}$ because $P_{n}$ is a rank-one matrix, making it easier to handle. Fortunately, the Weingarten calculus for $P_{n}=(p_{ij})_{1\leq i,j\leq n}$ has been settled in a nice form. In fact, the moment for the form
@@
is explicitly given by a simple rational function in $n$. We provide many examples of calculating the moments of matrix elements of $R_{n}$ and $P_{n}$. For example, we obtain
@@
if $k<n$. From the construction of the virtual isometry, the moments for $P_{n}$ essentially contain all the information of a Haar-distributed unitary matrix $U_{n}$. Therefore, this allows for the calculation of the moments of Haar-distributed unitary matrices in a way that is different from Weingarten calculus theoretically.
Our second main result is Theorem 3.3, in which we give an explicit relationship is established between the unitary Weingarten functions of dimensions $n$ and n1n-1. More especially, if $n>k$, the Weingarten function $\operatorname{Wg}_{k,n}$ on the symmetric group $S_{k}$ is expressed as a convolution product in the group algebra of $S_{k}$:
@@
where $\operatorname{as}_{k,n}$ is a new class function on $S_{k}$ and $\operatorname{Wg}_{k,n-1}$ is the Weingarten function of lower dimension n1n-1. We named the function $\operatorname{as}_{k,n}$ an ascension function and investigated its basic properties. For example, finding the inverse of function $\operatorname{as}_{k,n}$ with respect to the convolution product in the symmetric group algebra is straightforward. The assumption $n>k$ is necessary in the above formula. Although a similar expression can be derived for the case when $n=k$ with minor adjustments, it is important to emphasize that the resulting expression does not apply universally for all values of $n$ and $k$. Consequently, we have obtained a new representation of the Weingarten function utilizing the ascension functions and related constructs (see Proposition 3.11).
@@
At the end of the introduction, we comment on symbols used. Throughout this article, we use the notation
@@
for a complex number $a$ and a positive integer $k$. For convenience, we also set $a^{\uparrow 0}=1$. For clarity, we will use both $\delta_{ij}$ and $\delta(i,j)$ as Kronecker’s delta. The symbol $I_{n}$ represents the identity matrix of size $n$.
@@ Acknowledgments
B. C. is supported by JSPS Grant-in-Aid Scientific Research (A) no. 25H00593, and Challenging Research (Exploratory) no. 20K20882 and 23K17299.
S. M. is supported by JSPS Grant-in-Aid Scientific Research (C) no. 22K03233.
The idea to work on this project arose from discussions of B. C. with Sasha Bufetov and a talk by Ashkan Nikeghbali at the June 2023 RIMS conference on “Random Matrices and Applications”. We acknowledge discussions with them and the support of RIMS.
@@ 2.1. Complex reflection in the construction of virtual isometry
The following lemma is the key to our approach. For a proof, we refer to [BNN13, Proposition 3.1] or [BNR11, Theorem 1]. See also [BHNY08, Proposition 2.1].
@@
This lemma can also be obtained through a trick that is sometimes called the gaussianization method. Indeed, let $y_{1},\ldots,y_{n}$ be $n$ i.i.d standard complex gaussian variables. Then, the random vector $y/\|y\|_{2}$ has the same distribution as the uniformly distributed vector $x=(x_{1},\ldots,x_{n})$ on the norm $1$ sphere. Conversely, if $y$ and $x$ are independent, $\|y\|_{2}\cdot x$ has the same distribution as $y$. Using the complex version of Wick’s theorem and the fact that $\|y\|$ has a complex Chi-square distribution, thus satisfying $\mathbb{E}(\|y\|_{2}^{2k})=n(n+1)\cdots(n+k-1)$, the proof follows. For applications and related materials, we refer to Theorem 6.1 of [BC08]. We also refer to [DE02], in which a Chi-square distribution also appears in a tridiagonalization of a related random matrix.
@@ 2.3. Proof of the theorem
Let us rewrite Theorem 2.4 in another form. For an $n\times n$ matrix $A=(a_{ij})_{1\leq i,j\leq n}$ and each $1\leq k\leq n$, we denote by $a_{k\bullet}$ the sum of all entries in the $k$-th row of $A$ and by $a_{\bullet k}$ the sum of all entries in the $k$-th column of $A$:
@@
It is immediate to see that Theorem 2.4 is equivalent to Theorem 2.6.
@@
Since we are integrating on the unit sphere, the interchange of the infinite sum and the expectation follows from Fubini’s theorem. Applying Lemma 2.2 to each term, it survives only if equalities $a_{\bullet k}+b_{k\bullet}=a_{k\bullet}+b_{\bullet k}$ hold for all $k=1,2,\dots,n-1$, and if $m=l$ holds. In this case, we also have $a_{\bullet n}+b_{n\bullet}=a_{n\bullet}+b_{\bullet n}$ and $c\equiv 0\pmod{2}$, which leads to the following equations:
@@
Furthermore, from Lemma 2.2 again, we have
@@ 2.4. Some examples for reflection matrices
By utilizing Theorem 2.4 or Theorem 2.6, we can theoretically compute the moments for $(r_{ij})$. However, when there are many diagonal elements $r_{ii}$, there are many factors of the form $(1-p_{ii})$, which complicates the calculations using the theorems. While it seems complex in general, some calculations can be performed to a certain extent as follows.
@@
Let $m,l,q$ be nonnegative integers, and consider four sequences
@@
where $\alpha_{k}:=\mathsf{m}_{k}(\bm{i}\sqcup\bm{j}^{\prime})=\mathsf{m}_{k}(\bm{j}\sqcup\bm{i}^{\prime})$ for all $k=1,2,\dots,n$. Here the last factor may be zero.
@@
By the binomial theorem with $r_{ss}^{q}=(1-p_{ss})^{q}$, we have
@@
Using a similar method to the proof of Theorem 2.4, we view the last sum as the Gauss hypergeometric series, and obtain
@@
This concludes the proof for the case when $s<n$.
@@
Expand $|r_{ss}|^{2}=1-p_{ss}-\overline{p_{ss}}+p_{ss}\overline{p_{ss}}$ and apply Theorem 2.4. We omit the detailed calculations. ∎
@@
Using these, we obtain the following:
@@
Assume that the components of $\bm{i}$ are distinct and that $\bm{i}$ does not contain $n$. Then the moment
@@
We may suppose that $i_{h}=j_{h}$ for $1\leq h\leq f$ and $i_{h}\not=j_{h}$ for $f+1\leq h\leq m$. We define:
@@
The sequence $\bm{f}$ can be identified with the set $\{i_{1},i_{2},\dots,i_{f}\}$ since its components are distinct. We then have:
@@
where $\bm{k}$ runs over all subsets of $\bm{f}$. Applying Theorem 2.4 to each term, we find that a term corresponding to $\bm{k}$ survives only if $\bm{k}\sqcup\bm{j}_{0}$ is a permutation of $\bm{k}\sqcup\bm{i}_{0}$. In other words, $\bm{j}$ must be a permutation of $\bm{i}$. Furthermore, we can continue the calculation as follows:
@@
The following example plays a very important role from a theoretical point of view in Section 3.
@@
We can express the function given in (2.6) by using Kummer’s confluent hypergeometric function

### v3

@@ Abstract.
In this paper, we develop a novel approach to the Weingarten calculus by employing the notion of virtual isometries. Traditionally, Weingarten calculus provides explicit formulas for integrating polynomial functions over compact matrix groups with respect to the Haar measure, yet it faces limitations when evaluating high-degree integrals due to the non-invertibility of the associated matrices. We revisit these classical computations from a new perspective: by constructing Haar-distributed matrices as products of sequences of complex reflections, we derive new recursive structures for the Weingarten functions across different dimensions. This framework leads to two main results: (1) an explicit Weingarten calculus for complex reflections, yielding systematic moment computations for associated rank-one matrices, and (2) a novel convolution formula that connects Weingarten functions in dimension $n$ to those in dimension n1n-1, through the introduction of ascension functions in the symmetric group algebra. Our approach not only provides a unified treatment for unitary groups, but also sheds light on the algebraic and probabilistic aspects of high-degree integral computations. We present several examples and applications.
@@ 1.1. Weingarten calculus
An important case, which we focus on in this paper, is the unitary group $\mathrm{U}(n)$, which consists of all $n\times n$ unitary matrices. This case is significant because every compact matrix group is conjugate to a subgroup of the unitary group. Let $U=(u_{ij})_{1\leq i,j\leq n}$ be a Haar-distributed matrix from $\mathrm{U}(n)$. By properties of the Haar measure, when $k$ and $l$ are distinct nonnegative integers, the expectation of the following expression is zero:
@@
where $i_{p},j_{p},i_{p}^{\prime},j_{p}^{\prime}\in\{1,2,\dots,n\}$ and $p\geq 1$. According to the Weingarten formula [COL03], when $k=l$, this moment can be expressed as a sum over the symmetric group $S_{k}$:
@@
Here $\delta(p,q)$ is the Kronecker delta, and $\operatorname{Wg}_{k,n}$ is a class function on $S_{k}$ known as the (unitary) Weingarten function. See also [SAM80]. The Weingarten function has various representations; however, when $k\leq n$, we can immediately obtain the representation
@@
The Weingarten function has been the subject of extensive research. This includes expansions using characters of symmetric groups, expressions through Jucys–Murphy elements, and series expansions as generating functions for monotone factorizations of permutations [NOV10, MN13]. In addition, a relation between the Weingarten functions on $S_{k}$ and those on $S_{k-1}$ has been established [CM17].
@@ 1.2. The problem of high degree integrals
A common shortcoming of Weingarten calculus is that if one wants to evaluate an integral of degree strictly more than $n$ on the unitary or orthogonal group of dimension $n$, the Weingarten matrix is no longer invertible. For details, see, for example, [CMN22]. There are several approaches to address this problem. The first one is to take the pseudo-inverse of the Weingarten matrix when it is not invertible. It is important to note that while a Weingarten formula exists irrespective of the degree, the Weingarten function is not unique when the degree becomes larger than the dimension.
We are aware of two additional routes to compute polynomial integrals of high degree against the Haar measure, and describe them briefly in what follows. Both workarounds are successful from a theoretical point of view in the sense that they allow us to compute the integrals in principle. However, at this point, we are unable to extract any useful information for concrete applications. The first attempt can be found as a remark of a paper by Collins and Śniady [CŚ06], and the authors, along with Fukuda, have a joint working project to expand on this topic [CFM25]. Roughly, the idea is that any formal computation in the dimension, while in principle only well defined for a dimension that is large enough – and potentially larger than the dimension in which one wants to compute, doing the calculation formally, simplifying, and then evaluating in the target dimension will systematically give the correct result.
The second route was taken by Cioppa and Collins in the following unpublished article [CC13]. Here, the idea is to replace the indicator functions which are simple to compute but do not yield an orthogonal Gram matrix by much more complicated functions obtained from the Gelfand-Zetlin basis, i.e. exhibit matrix elements of the commutant of tensor representations of the unitary group.
The third approach is novel; it is actually the motivation for this paper. It comes from the following simple idea: from the point of view of differential geometry and measure theory, $\mathrm{U}(n)/\mathrm{U}(n-1)=S^{2n-1}$, and thus $\mathrm{U}(n)=S^{2n-1}\times\mathrm{U}(n-1)$ and this extends to the level of product of uniform measures. For a rigorous statement, we refer to Lemma 2.1. Therefore, we can understand the Haar measure at the level $n$ from the Haar measure at the level n1n-1 and the uniform measure on the sphere. But it turns out that the uniform measure on the sphere can be understood with arbitrary order directly thanks to a Gaussianization trick. This has been used in [BC08], and we also refer to the remark after Lemma 2.2 for more details.
@@ 1.3. Virtual isometries
Due to [BNN13] (see also [MNN20]), virtual isometries can be constructed inductively as follows: We equip $\mathbb{C}^{n}$ with the canonical scalar product: $\langle\bm{v},\bm{w}\rangle=\sum_{j=1}^{n}v_{j}\overline{w_{j}}$ for $\bm{v}=(v_{1},\dots,v_{n})^{\mathtt{T}}$, $\bm{w}=(w_{1},\dots,w_{n})^{\mathtt{T}}\in\mathbb{C}^{n}$. Here, $A^{\mathtt{T}}$ and $\bm{v}^{\mathtt{T}}$ represent the transpose of a matrix $A$ or a vector $\bm{v}$. Let $\bm{e}_{1},\dots,\bm{e}_{n}$ be the standard orthonormal basis of $\mathbb{C}^{n}$. Consider a sequence $(\bm{x}_{n})_{n\geq 1}$ of column vectors, where each $\bm{x}_{n}$ lies on the complex unit sphere of $\mathbb{C}^{n}$ for $n\geq 1$. We define a sequence of unitary matrices $r_{n}\in\mathrm{U}(n)$ in the following way:
@@
The virtual isometry is consistent with the Haar measure. Indeed, if $(X_{n})_{n\geq 1}$ is a sequence of independent random vectors, each uniformly distributed on the complex unit sphere of $\mathbb{C}^{n}$. If we denote by $R_{n}$ and $U_{n}$ the corresponding random matrices $r_{n}$ and $g_{n}$, respectively, in the above constructions, then $U_{n}$ follows the Haar measure on $\mathrm{U}(n)$.
A similar construction method for the Haar measure was also provided by A. Hurwitz (see [DF17]). Virtual isometries have proven to be valuable in the analytic aspects of random matrix theory, enabling the attainment of almost sure results due to the correlations they provide between sequences of Haar-distributed unitary matrices (see [ASS22, CNN17]).
@@ 1.4. Main results
The first main result is Theorem 2.4, which focuses on a random complex reflection $R_{n}$ used in a construction involving virtual isometries. We establish Weingarten calculus for $R_{n}$, which provides a systematic method for computing the moments of the matrix elements of $R_{n}$. Instead of working directly with $R_{n}$, we will deal with $P_{n}:=I_{n}-R_{n}$ because $P_{n}$ is a rank-one matrix, making it easier to handle. Fortunately, the Weingarten calculus for $P_{n}=(p_{ij})_{1\leq i,j\leq n}$ has been established in a concise form. In fact, the moment for the form
@@
is explicitly given by a simple rational function in $n$. We provide many examples of the calculation of the moments of matrix elements of $R_{n}$ and $P_{n}$. For example, we obtain
@@
if $k<n$. From the construction of the virtual isometry, the moments for $P_{n}$ essentially contain all the information of a Haar-distributed unitary matrix $U_{n}$. Therefore, this provides a theoretical method to calculate the moments of Haar-distributed unitary matrices that differs from the traditional Weingarten calculus.
Our second main result is Theorem 3.3, in which an explicit relationship is established between the unitary Weingarten functions of dimensions $n$ and n1n-1. More specifically, if $n>k$, the Weingarten function $\operatorname{Wg}_{k,n}$ on the symmetric group $S_{k}$ is expressed as a convolution product in the group algebra of $S_{k}$:
@@
where $\operatorname{as}_{k,n}$ is a new class function on $S_{k}$ and $\operatorname{Wg}_{k,n-1}$ is the Weingarten function of lower dimension n1n-1. We call the function $\operatorname{as}_{k,n}$ an ascension function and investigated its basic properties. For example, finding the inverse of the function $\operatorname{as}_{k,n}$ with respect to the convolution product in the symmetric group algebra is straightforward. The assumption $n>k$ is necessary in the above formula. Although a similar expression can be derived for the case when $n=k$ with minor adjustments, it is important to emphasize that the resulting expression does not apply universally for all values of $n$ and $k$. Consequently, we have obtained a new representation of the Weingarten function utilizing the ascension functions and related constructs (see Proposition 3.11).
@@
At the end of the introduction, we explain the notation used. Throughout this article, we use the notation
@@
where $a$ is a complex number and $k$ is a positive integer. For convenience, we also set $a^{\uparrow 0}=1$. For clarity, we will use both $\delta_{ij}$ and $\delta(i,j)$ to denote the Kronecker delta. The symbol $I_{n}$ represents the identity matrix of size $n$.
@@ Acknowledgments
B. C. is supported by JSPS Grant-in-Aid for Scientific Research (A) no. 25H00593, and Challenging Research (Exploratory) nos. 20K20882 and 23K17299.
S. M. is supported by JSPS Grant-in-Aid for Scientific Research (C) no. 22K03233.
The idea to work on this project arose from discussions between B. C. and Sasha Bufetov, and from a talk by Ashkan Nikeghbali at the June 2023 RIMS conference on “Random Matrices and Applications”. We are grateful for these stimulating conversations and for the support provided by RIMS.
@@ 2.1. Complex reflection in the construction of virtual isometry
The following lemma is the key to our approach. For a proof, we refer to [BNN13, Proposition 3.1] or [BNR11, Theorem 1]. See also [BHN+08, Proposition 2.1].
@@
This lemma can also be obtained through a trick that is sometimes called the gaussianization method. Indeed, let $y_{1},\ldots,y_{n}$ be $n$ i.i.d standard complex Gaussian variables. Then, the random vector $y/\|y\|$ has the same distribution as the uniformly distributed vector $x=(x_{1},\ldots,x_{n})$ on the complex unit sphere. Conversely, if $y$ and $x$ are independent, $\|y\|\cdot x$ has the same distribution as $y$. Using the complex version of Wick’s theorem and the fact that $\|y\|$ has a complex Chi-square distribution, thus satisfying $\mathbb{E}(\|y\|^{2k})=n(n+1)\cdots(n+k-1)$, the proof follows. For applications and related materials, we refer to Theorem 6.1 of [BC08]. We also refer to [DE02], in which a Chi-square distribution also appears in a tridiagonalization of a related random matrix.
@@ 2.3. Proof of the theorem
We rewrite Theorem 2.4 in an equivalent form. For an $n\times n$ matrix $A=(a_{ij})_{1\leq i,j\leq n}$ and for each $1\leq k\leq n$, we denote by $a_{k\bullet}$ the sum of all entries in the $k$-th row of $A$ and by $a_{\bullet k}$ the sum of all entries in the $k$-th column of $A$:
@@
It is straightforward to verify that Theorem 2.4 is equivalent to Theorem 2.6.
@@
Since we are integrating on the unit sphere, the interchange of the infinite sum and the expectation follows from Fubini’s theorem. Applying Lemma 2.2 to each term, a term survives only if the equalities $a_{\bullet k}+b_{k\bullet}=a_{k\bullet}+b_{\bullet k}$ hold for all $k=1,2,\dots,n-1$, and if $m=l$ holds. In this case, we also have $a_{\bullet n}+b_{n\bullet}=a_{n\bullet}+b_{\bullet n}$ and $c\equiv 0\pmod{2}$, which leads to the following equations:
@@
Furthermore, applying Lemma 2.2 again, we obtain
@@ 2.4. Some examples for reflection matrices
By utilizing Theorem 2.4 or Theorem 2.6, we can theoretically compute the moments for $(r_{ij})$. However, when many diagonal elements $r_{ii}$ appear, many factors of the form $(1-p_{ii})$ arise, which complicates the calculations using the theorems. While it seems complex in general, some calculations can be performed to some extent as follows.
@@
Let $m,l,q$ be nonnegative integers, and consider four sequences:
@@
where $\alpha_{k}:=\mathsf{m}_{k}(\bm{i}\sqcup\bm{j}^{\prime})=\mathsf{m}_{k}(\bm{j}\sqcup\bm{i}^{\prime})$ for all $k=1,2,\dots,n$. Note that the last factor may be zero.
@@
Applying the binomial theorem to $r_{ss}^{q}=(1-p_{ss})^{q}$, we have
@@
Using a similar method to that in the proof of Theorem 2.4, we recognize the last sum as a Gauss hypergeometric series and obtain
@@
This concludes the proof for the case $s<n$.
@@
Expand $|r_{ss}|^{2}=1-p_{ss}-\overline{p_{ss}}+p_{ss}\overline{p_{ss}}$ and apply Theorem 2.4. The detailed calculations are omitted. ∎
@@
Using these results, we obtain the following:
@@
Assume that the components of $\bm{i}$ are distinct and that none of them equals $n$. Then the moment
@@
We may assume that $i_{h}=j_{h}$ for $1\leq h\leq f$, and $i_{h}\not=j_{h}$ for $f+1\leq h\leq m$. We define:
@@
Since the components of $\bm{f}$ are distinct, the sequence $\bm{f}$ can be identified with the set $\{i_{1},i_{2},\dots,i_{f}\}$. We then have:
@@
where $\bm{k}$ runs over all subsets of the set $\bm{f}$. Applying Theorem 2.4 to each term, we find that a term corresponding to $\bm{k}$ is nonzero only if $\bm{k}\sqcup\bm{j}_{0}$ is a permutation of $\bm{k}\sqcup\bm{i}_{0}$. In other words, $\bm{j}$ must be a permutation of $\bm{i}$. Furthermore, the calculation continues as follows:
@@
The following example plays an important theoretical role in Section 3.
@@
We can express the function given in (2.6) using Kummer’s confluent hypergeometric function